40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1405     Entry time: Wed Apr 27 18:36:34 2011
Author: Alastair & Koji 
Type: Laser 
Category: GYRO 
Subject: Misalignment tests 

Since we are wondering what effect pointing may have (in particular with respect to coupling into the cavity and through faradays and onto PDs  etc)  we decided to do a little deliberate misalignment to see if the noise was impacted at all.

Firstly we misaligned the beam into the cavity.  We used only the horizontal actuator on the last mirror before the cavity for the CCW beam.  Looking at the trans PD DC voltage we moved from the initial max transmission (with the plastic cover in place over the 2 input mirrors) to lower transmission by moving the actuator CW and then CCW.

The result is that at lower frequency the noise was slightly improved with misalignment.  The misalignmend in the CW and CCW directions produced strangely similar results in that we get a factor of a couple at low freq.  At frequencies around 1Hz and above the noise gets worse due to the box being removed.  We confirmed this was the cause by going back to the max transmission (now only 740mV) with the box off and seeing that the higher value stayed at around 1Hz.  Interestingly after this moving around the noise at max transmission was now a little better than when we started.

Next we put the mirror back to the max transmission, and put foil over the small path between box and vacuum system which was worrying Koji.  You can see the effect most clearly in the second pdf below.  The noise between 0.4Hz and 6Hz is improved by a reasonable amount.

The final test we did was to misalign the beam onto the CCW PD by a small amount.  It would have been nice to quantify this and to know where to put the beam back to, but the DC output is so noisy (really terrible) that it is just impossible.  I misaligned it gradually using the horizontal actuator on the mirror right before the lens that focuses the beam onto the PD.  I moved it until the cavity would not lock, and then came back just a little until it would lock again.  It is much further toward the edge of the PD now, though it was difficult to tell whether any of the beam was clipping on thed edge (at least I couldn't see any through the viewer).  As you can see in the last pdf below the spectrum is quite a bit worse.

So while this doesn't give us the answer, it perhaps gives some clues as to what is going on here.

I finished up by trying to realign the beam on the CCW PD.  Since the DC level is messy it is difficult to know what to use for this.  I used the viewer, and then swept the cavity and got the error signal up on the scope.  It seemed like the beam was already giving the full error signal, and that trying to move it around using the mirror before the lens didn't give any difference  to the height of the error signal until the beam got right off the edge of the PD (around 7 full turns of the actuator in either direction).  One last FFT showed that the noise is back down to the level it was at before I misaligned the beam on the PD.

PS:  I also notice that we don't have the vacuum system pumped down at the moment.  Clearly it's not making a great deal of difference, but I guess we should pump it back down again soon.

Attachment 1: align_to_cavity.pdf  153 kB  | Hide | Hide all
align_to_cavity.pdf
Attachment 2: foil_over_air_gap.pdf  47 kB  | Hide | Hide all
foil_over_air_gap.pdf
Attachment 3: misalign_ccw_pd.pdf  88 kB  | Hide | Hide all
misalign_ccw_pd.pdf
ELOG V3.1.3-